OpenStudy (anonymous):

What is the general solution to? dy/dx = 2xy

OpenStudy (rational):

separate variables and integrate

OpenStudy (anonymous):

Do you mean to have it like In 2x dx In y dy

OpenStudy (rational):

\[\dfrac{dy}{dx} = 2xy\] \[\dfrac{1}{y}~dy = 2x ~dx\] \[\int \dfrac{1}{y}~ dy = \int 2x ~dx\]

OpenStudy (anonymous):

Okay then we need to simplify or add a + C

OpenStudy (anonymous):

Wow, thank you!

OpenStudy (rational):

\[\dfrac{dy}{dx} = 2xy\] \[\dfrac{1}{y}~dy = 2x ~dx\] \[\int \dfrac{1}{y}~ dy = \int 2x ~dx\] \[\ln |y| = x^2 + \ln C\] \[y = Ce^{x^2}\]

OpenStudy (rational):

corrected above^

OpenStudy (anonymous):

Ok, I see what you added. Thank you again

OpenStudy (rational):

you're welcome :D